OpenStudy (anonymous):

What polynomial has roots of -6, 1, and 4 ? x3 - 9x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 9x2 + 14x - 24

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Here's a handy tip: the value of the polynomial will be 0 at a root.

OpenStudy (whpalmer4):

Plug in x=-6 and evaluate. If you get something other than 0, move on to the next polynomial. If you have multiple polynomials that give you 0 for x = -6, try x=1, and then x = 4. Actually, probably easiest to do x = 1 first, because the arithmetic is easier!

OpenStudy (anonymous):

simplify (x+6)(x-1)(x-4)=0

OpenStudy (whpalmer4):

it's an okay assumption here, because the leading terms of all the choices have 1 as the coefficient, but it isn't as useful as a general tactic because it won't work if there was a constant factor as well. All of these equations have those roots, for example :-) \[2 x^3+2 x^2-52 x+48,\\3 x^3+3 x^2-78 x+72,\\4 x^3+4 x^2-104 x+96,\\5 x^3+5 x^2-130 x+120,\\6 x^3+6 x^2-156 x+144,\\7 x^3+7 x^2-182 x+168,\\8 x^3+8 x^2-208 x+192,\\9 x^3+9 x^2-234 x+216,\\10 x^3+10 x^2-260 x+240\]

OpenStudy (whpalmer4):

@prettigirlrock0804 how are we doing? Do you know about synthetic division and substitution yet?

OpenStudy (anonymous):

@whpalmer4 yes i know what a synthetic division is

OpenStudy (anonymous):

and sorry for not responding the other day my computer wasnt acting right

OpenStudy (whpalmer4):

How about now? Do you have an answer now? :-)

OpenStudy (anonymous):

hold on a sec let me work it out

OpenStudy (anonymous):

do you have to do the synthetic division for all of them

OpenStudy (whpalmer4):

You don't need to do any synthetic division here. You're trying to find the polynomial whose roots are -6, 1, and 4. Here are some approaches you can use: 1) build the polynomial in factored form which has roots -6, 1, 4, expand it (multiply it out), and compare with the answer choices. 2) evaluate each polynomial at x = -6, x = 1, x = 4 and make sure that all 3 produce a value of 0. If not, that polynomial does not have that set of roots. Now, for 2), using synthetic substitution (which is just like synthetic division) makes the process of evaluating the polynomials at each of those 3 values much faster than doing something like \[(-6)^3-9(-6)^2-22(-6)+24 = \] Do you need further explanation of either 1 or 2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

for the prob that you did when telling me about part 2 i got 264

OpenStudy (whpalmer4):

Right. So that means that that polynomial does not have a root at x = -6. You don't need to do any further testing of roots with that polynomial, as it is not a possible correct answer.

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

Actually, "right" only applies to the fact that you didn't get 0 for the result. I get -384. \[(-6)^3 -9(-6)^2-22(-6)+24 = -216-324+132+24 = -384\]

OpenStudy (anonymous):

so are you not suppose to get a zero

OpenStudy (whpalmer4):

Here's how you do it with synthetic substitution: -6 | 1 -9 -22 +24 ----------------------- 1 -6 | 1 -9 -22 +24 -6 ----------------------- 1 -15 -6 | 1 -9 -22 +24 -6 90 ----------------------- 1 -15 68 -6 | 1 -9 -22 +24 -6 90 -408 ----------------------- 1 -15 68 -384

OpenStudy (whpalmer4):

You ARE supposed to get a 0 for f(x) if x is a root of f(x)

OpenStudy (anonymous):

ok i got the final answer and its D

OpenStudy (whpalmer4):

For example, let's take the polynomial \[(x+1)(x-2) = x^2 -2x + 1x -2= x^2-x-2\] This has roots x = -1 and x = 2 If you plug in x = -1: \[(-1)^2-(-1)-2) = 1+1-2 = 0\]so x=-1 is a root If you plug in x = 1: \[(1)^2-(1)-2 = 1-1-2 = -2\]so x = 1 is not a root

OpenStudy (whpalmer4):

Really? What happens when you plug in x = 4? \[ x^3 + 9x^2 + 14x - 24\] \[(4)^3+9(4)^2+14(4)-24\] That's going to equal 0?!?

OpenStudy (anonymous):

yeah i just realized that

OpenStudy (anonymous):

ugh this is making me mad

OpenStudy (anonymous):

i tried all of them but yet cant find the correct answer

OpenStudy (anonymous):

@whpalmer4 hello are you still there

OpenStudy (whpalmer4):

Show me what you got when you evaluated the second polynomial

OpenStudy (whpalmer4):

\[x^3+x^2-26x+24\]

OpenStudy (anonymous):

-69

OpenStudy (whpalmer4):

-6 1 1 -26 24 -6 30 -24 --------------- 1 -5 4 0 so -6 is a root 1 1 1 -26 24 1 2 -24 -------------- 1 2 -24 0 so 1 is a root 4 1 1 -26 24 4 20 -24 ------------- 1 5 -6 0 so 4 is a root Or doing it the hard way: \[(-6)^3 + (-6)^2-26(-6)+24 = -216 + 36 +156 + 24 = 0\] \[(1)^3+(1)^2-26(1)+24 = 1+1-26+24 = 0\] \[(4)^3+(4)^2-26(4)+24 = 64+16-104+24=0\] Again, all 3 numbers are roots.

OpenStudy (whpalmer4):

I'm not sure which value of x you were using when you got -69, but you shouldn't get -69 for any of the 3 in that polynomial...if you show me your work, we can try to find the error so you don't make it or similar errors in the future.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x ^{3}-x ^{2}-26x+24\] \[-6^{3}-(-6)^{2}-26(-6)+24\]

OpenStudy (whpalmer4):

Okay, one problem is that we did different polynomials :-) It appears I told you to try the second one, but then I actually used the third one myself. In my defense, I did show the polynomial I was suggesting... Please try again with \[x^3+x^2-26+24\] Also, be very careful when writing powers of negative numbers. \[-6^3 = -(6^3)\]though you get the same answer either way. However, \[-6^2 \ne (-6)^2\]and in both cases, the one with parentheses is the correct one.

OpenStudy (anonymous):

i think i had messed up on my multiplication somewhere

OpenStudy (whpalmer4):

Get in the habit of always writing ( ) around a negative number when raising it to a power, because if the power is an even number, you'll probably get the wrong answer if you don't.

OpenStudy (whpalmer4):

So that should have been\[(-6)^3-(-6)^2-26(-6) + 24\]which evaluates to\[-216-36+156+24 = -72\]

OpenStudy (whpalmer4):

Sigh. I made another typo. Please try again with \[x^3+x^2-26x+24\]:-)

OpenStudy (anonymous):

yeah i found out where i have messed up

OpenStudy (anonymous):

and ok

OpenStudy (whpalmer4):

Did you see how my synthetic substitution thing made evaluation of the polynomials easier?

OpenStudy (anonymous):

with the -6